OpenStudy (anonymous):

f(x) = 4x2 + 3x + 5 . using the Mean Value Theorem on the interval [-1,1]. That is, find a number c at which the instantaneous rate of change of the function f equals the average rate of change of f over the interval [-1,1].

OpenStudy (anonymous):

i used the mean value theorm f'(c)=f(1)-f(-1)/1-(-1)

myininaya (myininaya):

and you can find f(1) and f(-1) from f(x)=4x^2+3x+5

OpenStudy (anonymous):

which equate to 8x+3=3

OpenStudy (anonymous):

does that mean no c value exists?

myininaya (myininaya):

nope that means you need to solve 8c+3=3 for c

OpenStudy (anonymous):

c=0/8

myininaya (myininaya):

if c occurs in the interval [-1,1] then there does exist a c in the interval [-1,1] such that f'(c)=(f(b)-f(a))/(b-a)

myininaya (myininaya):

c=0 is a number between -1 and 1

myininaya (myininaya):

you are done!

OpenStudy (anonymous):

oh right! thanks

myininaya (myininaya):

You do understand that 0/8=0 (and not undefined) if we had 8/0 (we wouldn't have found a number)

OpenStudy (anonymous):

yea...........i got confused about that i was thinking that it was undefined opps

myininaya (myininaya):

Yeah a lot of people do get confused about that always tried to tell them over and over if 0 is on top, you have 0 however if you have 0 on bottom, it is undefined

myininaya (myininaya):

but if you have study limits you also know 0/0 is indeterminate form

OpenStudy (anonymous):

thanks